OpenStudy (anonymous):

Find the distance (in metres) for speed 8×100000cm/s for 2 mins Full explanation pls

OpenStudy (wolfe8):

First you need to understand what speed is. Speed is the distance traveled in a certain time. You are given a speed in cm/s which means that the object travels a certain cm in 1 second. So here it travels 8 X 100000cm every second. You want for 2 minutes which have 2 X 60 seconds in them. What would your solution be?

OpenStudy (hesan):

Since S=Vt where S= distance , V = velocity and t= time we have to convert all valuesto SI system i.e v= 800000cm/s = 800000/100 = 8000 m /s t= 2 min = 2x 60 = 120 sec NOW S = Vt = (8000)(120) = 960000m This is the Complete poosible solution of this question May be this HElps u.

OpenStudy (hesan):

Where S= vt is the formula for horizontal motion of the body .

OpenStudy (anonymous):

Now I understood it thanks alot :)

OpenStudy (hesan):

Thanks